Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Ring Heater)
fafone, giri, rocchi - 23:04 Friday 19 February 2021 (50840) Print this report
SR Caliberation

At the beginning of the shift, we checked the illumination of HWS-DET and after confirming that it was ok (figure 1),Diego disabled the B5 drift control.
We started our acquisition and at 16.07 UTC, we switched on the SR RH with about 10W. The centre of the curvature seems to be in the same position (well-aligned) like yesterday (figure 2) (50832).
We also checked that the position of the beam is constant in time, thanks to the loop engaged by Paolo (50834).
We decided to keep the RH on for the whole night, so that the lens in the SR reaches the steady state. We will switch it off early in the morning.

Images attached to this report
Comments to this report:
fafone, rocchi - 12:49 Saturday 20 February 2021 (50842) Print this report

Since the RH curvature has not yet reached the steady state, we decided to keep the SR RH on for few more hours.

lumaca, nardecchia, rocchi - 21:52 Saturday 20 February 2021 (50844) Print this report

The SR RH has been switched OFF at 20:34 UTC.

Search Help
×

Warning

×